1) find the area of a parallelogram with height 6inches and base 11 inches. 2)find the perimeter and area of a square with a 12 ft. side. 3) find the area of a triangle with base 13ft. and height 6 feet. thank you so much for your help
Answer by checkley71(8403)

1) 6*11=66 IS THE AREA.
2) 4*12=48 IS THE PERIMETER.
3)13/2*6=6.5*6=39 AREA.